OpenStudy (anonymous):

(1-cos^2x)(1-tan^2x) = sin^x-2sin^4x/1-sin^2x prove ls=rs

OpenStudy (turingtest):

I think there is a typo in the rhs what is sin^x ?

OpenStudy (anonymous):

sin^2x

OpenStudy (turingtest):

\[(1-\cos^2x)(1-\tan^2x)={\sin^2x-2\sin^4x\over1-\sin^2x}\]right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I did left side, and so far I got sin^2x (1-sin^2x/cos^2x) is taht right?

OpenStudy (turingtest):

yeah that is correct, though I think it may be easier to work the right side first...

OpenStudy (anonymous):

I trued but didn't really know what to do, I got stuck

OpenStudy (anonymous):

*tried

OpenStudy (turingtest):

that 2 doesn't seem to belong... are you sure there is a coefficient of 2 on the right?

OpenStudy (anonymous):

yess

OpenStudy (turingtest):

\[(1-\cos^2x)(1-\tan^2x)=\sin^2x(1-{\sin^2x\over\cos^2x})=\sin^2x-{\sin^4x\over\cos^2x}\]oh yes, that is right now take cos^2x as a common denominator so you can put the fractions together, then use cos^2x=1-sin^2x

OpenStudy (turingtest):

\[(1-\cos^2x)(1-\tan^2x)=\sin^2x(1-{\sin^2x\over\cos^2x})=\sin^2x-{\sin^4x\over\cos^2x}\]\[={\sin^2x\cos^2x-\sin^4x\over\cos^2x}\]

OpenStudy (turingtest):

now like I say\[\cos^2x=1-\sin^2x\]and you're done

OpenStudy (anonymous):

wait, so you just made the top cos^2x and bottom cos^2x into 1-sin^2x and brought it to the bottom?

OpenStudy (anonymous):

and but it answer should come up to sin^2x-2sin^2x/1-sin^2x

OpenStudy (turingtest):

\[(1-\cos^2x)(1-\tan^2x)=\sin^2x(1-{\sin^2x\over\cos^2x})=\sin^2x-{\sin^4x\over\cos^2x}\]\[=\sin^2x\cdot{\cos^2x\over\cos^2x}-{\sin^4x\over\cos^2x}={\sin^2x\cos^2x-\sin^4x\over\cos^2x}\]now \[\cos^2x=1-\sin^2x\]and simplify and you do get that answer

OpenStudy (turingtest):

I multiplied the \(\sin^2x\) by\[\frac{\cos^2x}{\cos^2x}\]

OpenStudy (anonymous):

oh okayy, thank youu!

OpenStudy (turingtest):

welcome :)